OpenStudy (anonymous):

I need help converting this exponential equation "4^y=8" to a logarithmic equation so I can graph it. I know 4 is the base, y is the exponent and 8 is the argument. So the log expression is "log subscript 4 eight=y" right? My question now is, how do I graph it?

hero (hero):

log both sides then remember that \[\log(a^b) = a \log(b)\]

OpenStudy (anonymous):

log(4^y)= 4 log y? or is it supposed to be ?